# NN-ResDMD: Learning Koopman Representations for Complex Dynamics with Spectral Residuals

### v1

@@ A.8.1 Visualization and Clustering Performance
By applying MDS, the high-dimensional eigenfunction-based features are reduced to a low-dimensional space. For illustration, we present the results of reducing the feature space to two dimensions (Figure 6B-E). The NN-ResDMD reduced features for the six types of trials (corresponding to the six video stimuli) are well-separated for all five mice (Figure 6B). In contrast, the Hankel-DMD features show no clear clustering structure (Figure 6C). Similarly, the features produced by EDMD with an RBF basis and Kernel ResDMD do not show clear separability (Figure 6D-E, Appendix Figure 9B-D, Appendix Figure 10B-D).
Figure 8: State Partition performance of eigenfunctions for NN-ResDMD and Hankel-DMD in 2D space visualized with UMAP (A) and t-SNE (B).
@@
Figure 11: Davies-Bouldin Indices evaluating the clustering performance of dynamical components learned by four methods (NN-ResDMD, Hankel DMD, EDMD+RBF, and Kernel ResDMD) across five mice. Comparisons are shown using UMAP (A) and t-SNE (B).
A.8.2 Clustering Quality Metrics
We further quantified the clustering quality by calculating the Davies-Bouldin Index (DBI) for both Koopman decomposition methods across all mice (Figure 6F). The DBI is designed to assess the compactness of clusters and the separability between them. A lower DBI indicates better clustering performance. NN-ResDMD features yield significantly lower DBI scores compared to other methods, confirming that NN-ResDMD produces more clearly defined clusters corresponding to the ground truth trials. Similar clustering results are observed with UMAP and t-SNE (see Appendix Figure 11), further supporting the superior performance of NN-ResDMD in capturing the latent dynamic structure compared to the other classical methods.
A.9 Source code
For reproducibility, the source code will be available at the following anonymous URL: https://anonymous.4open.science/r/ICLR-7305-PROJ. A full version of the codebase will be released upon acceptance of the paper.

### v3

@@ A.10.1 Visualization and clustering performance
By applying MDS, the high-dimensional eigenfunction-based features are reduced to a low-dimensional space. For illustration, we present the results of reducing the feature space to two dimensions (Figure 6B-E). The ResKoopNet reduced features for the six types of trials (corresponding to the six video stimuli) are well-separated for all five mice (Figure 6B). In contrast, the Hankel-DMD features show no clear clustering structure (Figure 6C). Similarly, the features produced by EDMD with an RBF basis and Kernel ResDMD do not show clear separability (Figure 6D-E, Appendix Figure 14B-D, Appendix Figure 15B-D).
Figure 10: Trial-averaged neural responses to 6 video stimuli (marked by 6 states).
Figure 11: (A) Mean firing rates (calcium spike) of each mouse in response to 6 video stimuli (marked as states). (B) Clustering performance based on the mean firing rates of each mouse to 6 video stimuli, visualized with Multi-dimensional Scaling (MDS). (C) Same as (B) but visualized with UMAP. (D) Same as (B) but visualized with t-SNE
Figure 12: Comparison of clustering quality based on Davies-Bouldin Indices for clustering with mean firing rates and Koopman eigenfunctions learned from ResKoopNet, visualized in 2D space with Multi-dimensional Scaling (A), UMAP (B) and t-SNE (C).
Figure 13: State Partition performance with eigenfunctions for ResKoopNet (50 bases) and Hankel-DMD in 2D space visualized with UMAP (A) and t-SNE (B).
@@
Figure 16: Davies-Bouldin Indices evaluating the clustering performance of dynamical components learned by four methods (ResKoopNet(50 bases), Hankel DMD, EDMD+RBF, and Kernel ResDMD) across five mice. Comparisons are shown using UMAP (A) and t-SNE (B).
Figure 17: Davies-Bouldin Indices for different neural network hyperparameters in training ResKoopNet with 50 bases in an example mouse. Note that lower values indicate better clustering performance. With small layer size and layer number, the clustering performance is not stable but becomes robust when the layer size reached 200 and layer number reaches 3.
A.10.2 Clustering quality metrics
We further quantified the clustering quality by calculating the Davies-Bouldin Index (DBI) for both Koopman decomposition methods across all mice (Figure 6F). The DBI is designed to assess the compactness of clusters and the separability between them. A lower DBI indicates better clustering performance. ResKoopNet features yield significantly lower DBI scores compared to other methods, confirming that ResKoopNet produces more clearly defined clusters corresponding to the ground truth trials. Similar clustering results are observed with UMAP and t-SNE (see Appendix Figure 16), further supporting the superior performance of ResKoopNet in capturing the latent dynamic structure compared to the other classical methods.